Dsbubduhsdgywgd cgucgdugsudhwydgywguwdgygdwydgywdg
 
Value of................
 
Sin [1/2 cot^-1 (-3/4)]=

Grade:12

1 Answers

Sayam
6 years ago
=sin [1/2 cot-1(-3/4)]
=sin [1/2( pie-cot-1(3/4)]   =sin [1/2( pie-sin-1(4/5)]  
=sin [( pie/2-sin-1(4/5)/2]  =cos(sin-1(4/5)/2]    =cos(cos-12/root 5)
=2/root 5  …..Hope this clears your doubt.